# Adversarial KA

### v1

@@ 2. Background and statements
and repeat the construction of $g$ from the proof of Lemma 2.1. ∎
After this lemma, construct the universal tuples $\phi_{1},\ldots,\phi^{n}$ and approximations for each $f$ as in Lemma 3.2. Finally, apply the obtained analogue of Lemma 3.2 to fuel a non-linear recursion and thus prove Theorem 2.2.
@@
Question 3.4.
Do the analogues of Theorems 2.1, 2.2 hold when the set of adversaries is the entire homeomorphism group of $\mathbb{R}^{2n+1}$?
Question 3.5.
Do the analogues of Theorems 2.1, 2.2 hold for adversaries acting after the summation, as diagramed in Figure 2, for any interesting class of adversaries beyond the cases of: 1. Countable subsets of the affine group, and 2. The continuous group of translations $\{T\}$ as discussed in the final paragraph of §4.Conclusions (below)? Both countable and continuous families of $C^{2}$-diffeomorphisms look like an interesting case to consider.
@@ 4. Conclusions
All proofs of KA using the Baire category theorem, including ours, easily pass from a countable dense collection $\{f\}$ of functions for which a <<working>> $\phi$ has been found to all functions but taking a limit. It may therefore come as a surprise that we are not able to similarly take a limit to pass from a countable collection $\mathcal{H}$ of adversaries to its closure. This would be the hoped-for conclusion. What goes wrong? The problem is that our rule for producing the outer function $g^{f,h}$ depends delicately on the location of the <<forced>> values on the real line. The function $g^{f,h}$ at these values (both in the approximation step and after convergence) bounces around between the various $\frac{1}{3}f(\ell_{j,r})$, which initially means within $[\,-\frac{1}{3},\frac{1}{3}\,]$. If such large transitions occur at near and nearer <<forced>> values the Lipshitz constant of $g^{f,h}$ will diverge. As the adversary $h$ varies (think of rotating the plane and projecting) the <<forced>> values may cross each other. This means that the functions $\{g^{f,h}\}$ are not necessarily equi-continuous w.r.t. $h$.
To give a very simple instance of this problem consider an adversary which acts on $\mathbb{R}^{5}$ merely by Euclidian translation. So, $\mathbb{R}^{5}=\{T\}$ is (also) our group of allowed adversaries. If one translates any given coordinate axis, the <<forced>> values will not cross and, trivially, the adversary may be defeated by simply pre-composing $\{g^{T}_{1},\ldots,g^{T}_{5}\}$ by the inverse translation $T^{-1}$ — if we are allowed five independent outer functions, one for each hidden layer neuron. If we continue, though, to maintain the initial rules of our game, and ask for a single outer function $g^{T}$, then the independent axial translations of the <<forced>> values, now projected onto a single Real axis, will in general cross, causing a divergence of $\mathop{\text{Lip}}(g^{T})$ and a loss of equi-continuity. This is the simplest context in which we do not know how to answer our question. As explained in our discussion of affine adversaries, translations $\{T\}$ it makes no difference whether the adversarial action occurs before or after summation is taken at the hidden layer.

### v3

@@ 2. Background and statements
assume a constant value which we simply denote as $\Phi^{\{h^{i}\}}_{j,r}$.
We must further lower bound $N$ by requiring (from the uniform continuity of the function $f$) that $\left\lvert f(x)-f(x^{\prime})\right\rvert<\alpha$ for any tuples $x,x^{\prime}\in I^{n}$ located in the same $R_{j,r}$.
For each red $n$-rectangular solids $R_{j,r}$ choose some point $\ell_{j,r}\in R_{j,r}$.
Finally, define a piecewise linear $g^{f,\{h^{i}\}}\colon\mathbb{R}\to\mathbb{R}$ by the ‘forced’ values $g^{f,\{h^{i}\}}(\Phi^{\{h^{i}\}}_{j,r}):=\frac{1}{n+1}f(\ell_{j,r})$ for each red $n$-rectangular solids $R_{j,r}$. This definition makes sense since all $\Phi_{j,r}^{\{h^{i}\}}$ are distinct. Indeed, on two different red $n$-rectangular solids $R_{j,r}$ and $R_{j^{\prime},r^{\prime}}$ we have
$\Phi_{j,r}^{\{h^{i}\}}=\sum_{i=1}^{n}q_{i}\gamma_{i}\quad\text{and}\quad\Phi_{j^{\prime},r^{\prime}}^{h}=\sum_{i=1}^{n}q^{\prime}_{i}\gamma_{i},$
where $q_{i},q^{\prime}_{i}\in\mathbb{Q}$ and at least one of the inequalities $q_{i}\neq q_{i}^{\prime}$ holds, and that is why $\Phi_{j,r}^{\{h^{i}\}}\neq\Phi_{j^{\prime},r^{\prime}}^{\{h^{i}\}}$.
Again, extend $g^{f,h}$ from its ‘forced’ values to ensure that $\left\lVert g^{f,h}\right\rVert\leqslant\frac{1}{n+1}$.
$\operatorname{Approx}(g,\{h^{i}\})=\sum_{j=1}^{2n+1}g\circ\Phi_{j}^{\{h^{i}\}},$
we need to prove that
$\left\lvert f(x)-\sum_{j=1}^{2n+1}g^{f,\{h^{i}\}}\bigl(\Phi_{j}^{\{h^{i}\}}(x)\bigr)\right\rvert<\alpha+\frac{n}{n+1}\quad\text{for any}\quad x\in I^{n}.$
Take any $x\in I^{n}$. As we already know, there exist at least $n+1$ red $n$-rectangular solids $R_{j_{s},r_{s}}$, $s\in\{1,\ldots,n+1\}$, containing the point $x$. For them we have $g^{f,\{h^{i}\}}(\Phi_{j_{s},r_{s}}^{\{h^{i}\}})=\frac{1}{n+1}f(\ell_{j_{s},r_{s}})$. These values differ from $\frac{1}{n+1}f(x)$ by less than $\frac{\alpha}{n+1}$. Hence
$\left\lvert\sum_{s=1}^{n+1}g^{f,\{h^{i}\}}\left(\Phi^{\{h^{i}\}}_{j_{s}}(x)\right)-f(x)\right\rvert<\alpha.$
The other $n$ values $g^{f,\{h^{i}\}}\left(\Phi^{\{h^{i}\}}_{j}(x)\right)$ do not exceed $\frac{1}{n+1}$ by the absolute value. ∎
Lemma 4.2.
Let $n>1$ be an integer and $\lambda\in(\frac{2n+1}{2n+2},1)$. Let $\mathcal{H}=\{h_{q}\}$ be a countable set of homeomorphisms $\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$. There exist fixed $\phi^{1},\ldots,\phi^{n}\colon I\to I^{2n+1}$ such that given any $f\colon I^{n}\to\mathbb{R}$ and any $h^{1},\ldots,h^{n}\in\mathcal{H}$, there is a uniformly continuous $g\colon\mathbb{R}\to\mathbb{R}$ with $\left\lVert g\right\rVert\leqslant\frac{1}{n+1}\left\lVert f\right\rVert$ satisfying
$\left\lVert f-\operatorname{Approx}(g,\{h^{i}\})\right\rVert<\lambda\left\lVert f\right\rVert.$
Proof.
W.l.o.g. assume $\left\lVert f\right\rVert=1$ and let $f_{1},f_{2},\ldots$ be an infinite sequence of functions in $C(I^{n},\mathbb{R})$ all of norm $1$ and dense in the unit sphere of $C(I^{n},\mathbb{R})$. Take $\alpha:=\lambda-\frac{2n+1}{2n+2}$. The space $C(I,I^{2n+1})^{n}$ is a separable complete metric space, so by Lemma 4.1 and the Baire category theorem,
$\bigcap_{1\leqslant p,q<\infty}U_{f_{p},\{h^{i}\}_{q}}\neq\varnothing.$
(In fact, this intersection is a dense set of ‘2nd category’.) Choose $\phi=(\phi^{1},\ldots,\phi^{n})$ from this intersection. For this $\phi$ choose maps $g^{f_{p},\{h^{i}\}_{q}}$ from the definition of $U_{f_{p},\{h^{i}\}_{q}}$.
Now fix any $f\colon I^{n}\to\mathbb{R}$ and any $\{h^{i}\}=\{h^{i}\}_{q}$. We need to choose the appropriate $g$. For this, choose $f_{p}$ so that $\left\lVert f-f_{p}\right\rVert<\frac{1}{2n+2}$.
Now $g:=g^{f_{p},\{h^{i}\}_{q}}$ satisfies the requirements of the lemma, since
$\left\lVert f-\operatorname{Approx}(g,\{h^{i}\})\right\rVert\leqslant\left\lVert f-f_{p}\right\rVert+\left\lVert f_{p}-\operatorname{Approx}(g,\{h^{i}\})\right\rVert<\frac{1}{2n+2}+\alpha+\frac{n}{n+1}=\lambda.$
Proof of Theorem 2.2.
Take any $\lambda\in(\frac{2n+1}{2n+2},1)$. Take $\phi$ from Lemma 4.2. Fix any $f\colon I^{n}\to\mathbb{R}$ and $h^{1},\ldots,h^{n}\in\mathcal{H}$.
Let $g_{0}:=0$ and let $g_{m+1}$ be obtained by applying Lemma 4.2 for fixed $\phi$ to $f-\sum\limits_{k=0}^{m}\operatorname{Approx}(g_{k},\{h^{i}\})$ and $\{h^{i}\}$. Hence for any integer $m>0$
$\left\lVert f-\sum\limits_{k=0}^{m}\operatorname{Approx}(g_{k},\{h^{i}\})\right\rVert<\lambda\left\lVert f-\sum\limits_{k=0}^{m-1}\operatorname{Approx}(g_{k},\{h^{i}\})\right\rVert<\ldots<\lambda^{m}\left\lVert f\right\rVert.$
Since $\left\lVert g_{m}\right\rVert\leqslant\frac{1}{n+1}\left\lVert f-\sum\limits_{k=0}^{m-1}\operatorname{Approx}(g_{k},\{h^{i}\})\right\rVert\leqslant\frac{1}{n+1}\lambda^{m-1}\left\lVert f\right\rVert$, the functional series $\sum\limits_{m=0}^{\infty}g_{m}$ converges uniformly on $\mathbb{R}$ to a uniformly continuous function $g\colon\mathbb{R}\to\mathbb{R}$. Then
$\sum_{m=0}^{\infty}\operatorname{Approx}(g_{m},\{h^{i}\})=\operatorname{Approx}\left(\sum_{m=0}^{\infty}g_{m},\{h^{i}\}\right)=\operatorname{Approx}(g,\{h^{i}\}).$
Hence $\left\lVert f-\sum\limits_{m=0}^{\infty}\operatorname{Approx}(g_{m},\{h^{i}\})\right\rVert=0$, and so $f=\operatorname{Approx}(g,\{h^{i}\})$. ∎
5. Related work and positioning
See the overview [18] for the history of KA, its discussion and applications. The KA was well studied in the spirit of strengthening the restrictions imposed on the functions:
  • in [17], it was shown that KA fails if all the functions are required to be continuously differentiable,
  • in [5], it was shown that all the inner layer functions $\phi$ could be made 11-Lipschitz continuous,
  • it is known [14, 9] that all the inner layer functions $\phi$ could be made strictly increasing.
Note that in [3] the first constructive proof of KA theorem was given.
Instability of KA representations was known qualitatively. Our paper is probably the first to investigate the uniform stability of KA under the effects of collections of re-parameterizations, acting on the hidden layer.
@@ 5. Related work and positioning
Question 5.1.
Do the analogues of Theorems 2.1, 2.2 hold when the collection of re-parameterizations is the entire homeomorphism group of $\mathbb{R}^{2n+1}$?
Question 5.2.
Do the analogues of Theorems 2.1, 2.2 hold for re-parameterizations acting after the summation, as diagramed in Figure 2, for any interesting class of re-parameterizations beyond the cases of: 1. Countable subsets of the affine group, and 2. The continuous group of translations $\{T\}$ as discussed in the final paragraph of §6.Conclusions (below)? Both countable and continuous families of $C^{2}$-diffeomorphisms look like an interesting case to consider.
@@ 6. Conclusions
All proofs of KA using the Baire category theorem, including ours, easily pass from a countable dense collection $\{f\}$ of functions for which a ‘working’ $\phi$ has been found to all functions but taking a limit. It may therefore come as a surprise that we are not able to similarly take a limit to pass from a countable collection $\mathcal{H}$ of re-parameterizations to its closure. This would be the hoped-for conclusion. What goes wrong? The problem is that our rule for producing the outer function $g^{f,h}$ depends delicately on the location of the ‘forced’ values on the real line. The function $g^{f,h}$ at these values (both in the approximation step and after convergence) bounces around between the various $\frac{1}{3}f(\ell_{j,r})$, which initially means within $\frac{1}{3},\frac{1}{3}\,]$. If such large transitions occur at near and nearer ‘forced’ values the Lipshitz constant of $g^{f,h}$ will diverge. As the re-parameterization $h$ varies (think of rotating the plane and projecting) the ‘forced’ values may cross each other. This means that the functions $\{g^{f,h}\}$ are not necessarily equi-continuous w.r.t. $h$, and this turns out to be the main obstruction for the step from countable collections of homeomorphisms to continuous ones.
To give a very simple instance of this problem, consider a re-parameterization that acts on $\mathbb{R}^{5}$ merely by Euclidian translation. So, $\mathbb{R}^{5}=\{T\}$ is (also) our group of allowed re-parameterizations. If one translates any given coordinate axis, the ‘forced’ values will not cross and, trivially, the re-parameterization may be defeated by simply pre-composing $\{g^{T}_{1},\ldots,g^{T}_{5}\}$ by the inverse translation $T^{-1}$if we are allowed five independent outer functions, one for each hidden layer neuron. If we continue, though, to maintain the initial rules of our game, and ask for a single outer function $g^{T}$, then the independent axial translations of the ‘forced’ values, now projected onto a single Real axis, will in general cross, causing a divergence of $\mathop{\text{Lip}}(g^{T})$ and a loss of equi-continuity. This is the simplest context in which we do not know how to answer our question. As explained in our discussion of affine re-parameterizations, translations $\{T\}$ it makes no difference whether the re-parameterization action occurs before or after summation is taken at the hidden layer.
There is another natural question concerning the computability of all functions. We do not give the precise definition of a computable function $\mathbb{R}\to\mathbb{R}$ which can be found, for example, in [2, 20]. Analogously to [8, 2], we pose the following final question.
Question 6.1.
Does the class of functions $I^{n}\to\mathbb{R}$, implementable by neural networks with a single hidden layer and computable activation functions, computable weights, and a computable adversary acting as in Figure 1, coincide with the class of all computable functions $I^{n}\to\mathbb{R}$?